### v1

@@ Abstract
High-precision modeling of systems is one of the main areas of industrial data analysis. Models of systems, their digital twins, are used to predict their behavior under various conditions. We have developed several models of a storage system using machine learning-based generative models. The system consists of several components: hard disk drive (HDD) and solid-state drive (SSD) storage pools with different RAID schemes and cache. Each storage component is represented by a probabilistic model that describes the probability distribution of the component performance in terms of IOPS and latency, depending on their configuration and external data load parameters. The results of the experiments demonstrate the errors of 4–10 % for IOPS and 3–16 % for latency predictions depending on the components and models of the system. The predictions show up to 0.99 Pearson correlation with Little’s law, which can be used for unsupervised reliability checks of the models. In addition, we present novel data sets that can be used for benchmarking regression algorithms, conditional generative models, and uncertainty estimation methods in machine learning.
@@
Introduction Data storage systems (DSS) are of vital importance in the modern world. The emergence of big data has necessitated the development of storage systems with a larger capacity and lower costs to efficiently store and process the vast amounts of data generated [1, 2]. Performance modeling is essential in the development of such systems. The modeling allows engineers to analyze the system’s behavior under different conditions and helps them to find the optimal design of the system. It can be used for marketing purposes to estimate the system performance under the customer’s requirements. One more application is diagnostics and predictive maintenance, where model predictions are compared with actual measurements to detect failures and anomalies.
The key components of a typical DSS are controllers, fast cache memory, and storage pools. All data are stored in the pools consisting of several hard disk drives (HDDs) or solid-state drives (SSDs) united with raid schemes. The cache speeds up read and write operations with the most popular data blocks. And the controllers are responsible for data management and provide computational resources to process all input and output requests from users.
An early attempt to model SSDs was described in [3]. The authors argue that HDD performance models cannot be used for SSD due to certain unique SSD characteristics, such as low latency, slow update, and expensive block-level erasure. The black-box approach to model SSD performance is suggested as it requires minimal a priori information about the device. The researchers found that although black-box simulation does not produce high-quality predictions for HDDs, it can produce accurate ones for SSDs.
@@
Machine learning techniques have proved to be useful in the task of analyzing the disks. As an example, the main approaches discussed in [6] are based on machine learning. The paper uses a methodology that extracts low-cost history-aware features to train SSD performance models which predict response times to requests. The paper also utilizes space-efficient data structures such as exponentially decaying counters to track past activity with $O(1)$ memory and processing cost. Lastly, the paper uses machine learning models such as decision trees, ensemble methods, and feedforward neural networks to predict the storage hardware response time. All of these approaches are designed to make accurate SSD simulators more accessible and easier to use in real-world online scenarios.
@@
Another open-source simulator was developed in [8]. The authors introduce SimpleSSD, a high-fidelity holistic SSD simulator. Aside from modeling a complete storage stack, this simulator decouples SSD parallelism from other flash firmware modules, which helps to achieve a better simulation structure. SimpleSSD processes I/O requests through layers of a flash memory controller. The requests are then serviced by the parallelism allocation layer, which abstracts the physical layout of interconnection buses and flash disks. While being aware of internal parallelism and intrinsic flash latency, the simulator can capture close interactions among the firmware, controller, and architecture.
The idea of SimpleSSD has been further developed in [9]. SimpleSSD 2.0, namely Amber, was introduced to run in a full system environment. Being an SSD simulation framework, Amber models embedded CPU cores, DRAMs, and various flash technologies by emulating data transfers. Amber applies parallelism-aware readahead and partial data update schemes to mimic the characteristics of real disks.
@@
Although simulating SSD devices might have seemed an exciting research topic, older kinds of storage system has also received the scientific attention it deserves. One of the papers presents a series of HDD simulations to investigate the stiffness of the interface between the head and the disk. In [11], the authors develop a finite-element model of an HDD using a commercially available three-dimensional modeling package. The results of the numerical analysis gave insights into the characteristics of each type of shock applied to an HDD model.
@@
The evaluation of SSD performance has also been investigated in [14]. SSDs are built using an integrated circuit as memory, so manufacturers are interested in I/O traces of applications running on their disks. This paper proposes a framework for the accurate estimation of I/O trace execution time on SSDs.
@@
All of these solutions have a range of limitations. The first one is that they are focused mostly on the modeling of one single device such as HDD or SSD disks. The second is that most of them are based on the simulation of physical processes and the firmware stack inside the devices which require more resources for development. But one of the main disadvantages is that the models do not take into account vendor specifics of the devices due to the trade secrets. In result, errors of the performance simulation are quite high for practical use. Finally, HDDs and SSDs are part of DSS with unique architectures and software that influence the device performance and are not taken into account in the solutions described above.
@@
In this work, we present a data-driven approach for DSS modeling which does not have these limitations. All vendor specifics, architecture features, and software impact are learned directly from real performance measurements of the system. We present the performance modeling and analysis of a data storage cache, SSD, and HDD pools under random and sequential data loads. We provide novel data sets of IOPS and latency measurements for the system’s components for various configurations and load parameters. We consider two different approaches to learn distributions of the performance values, based on parametric and nonparametric generative models. We analyze the quality of the model’s predictions and compare them with a naive baseline algorithm. We also demonstrate a physics-inspired method for checking the reliability of the model predictions. The data sets provided in this work can be used for benchmarking regression algorithms, conditional generative models, and uncertainty estimation methods in machine learning.
Problem statement
Table 1: Input and output features of the simulation models
Model
Inputs
Outputs
Storage pool
@@
Load type, IO type, read fraction, block size, number of jobs, queue depth, number of disks, number of data, and parity blocks in RAID
IOPS, Latency
Storage cache
@@
Load type, IO type, read fraction, block size, number of jobs, queue depth
IOPS, Latency
@@
In this study, we consider the simulation of HDD and SSD storage pools, and cache, which are the key parts of data storage systems. The goal of this work is to predict the performance of these components for a given configuration and data load parameters. We describe the performance by the number of input and output operations per second (IOPS) and the average of their latencies. The data load is parameterized by: load type, IO type, read fraction, block size, number of jobs, and queue depth for each of these jobs. There are two load types we consider in the work: random and sequential. Each data load consists of a mixture of read and write operations. The IO type indicates whether we consider read or write operations in our simulations. The read fraction is the ratio of the number of read operations to the total number of all operations in the data load. Each operation processes one data block of a given size. And the data loads are generated by several jobs with a given queue depth.
@@
HDD and SSD storage pools also have configuration parameters. They are the total number of disks in a pool and the RAID scheme, which is described by the number of data and parity blocks. In this work, the cache has only one configuration, and we do not describe it by any parameter. The lists of input and output features in our simulation models are shown in Table 1.
@@
Let us denote the vector of input features as $x$, and the vector of outputs as $y$. Also suppose we have $n$ pairs of measurements $\{x_{i},y_{i}\}_{i=1}^{n}$ of output vectors $y_{i}$ for the given inputs $x_{i}$. These pairs are obtained by generating various data loads on a real data storage system and measuring the performances of its pool or cache in terms of IOPS and latency. The goal of this work is to estimate the conditional probability distribution $p(y|x)$ for the storage component using machine learning. Then, we predict the performance $\hat{y}_{j}$ for the unknown inputs $x_{j}$ by sampling from the learned distribution:
@@
Table 2: Data load parameters and their value ranges for the storage cache data set
Parameter
Value range
Load type
random
Block size
4, 8, 16, 32, 64, 128, 256 KB
Read fraction
0 - 100%
Number of jobs
1 - 64
Queue depth
1 - 16
@@
Table 3: Data load parameters and their value ranges for the storage SSD pool data set; random data loads
Parameter
Value range
Load type
random
Block size
4, 8, 16, 32, 64 KB
Read fraction
0 - 100%
Number of jobs
1 - 32
Queue depth
1 - 32
RAID (K+M)
1+1, 2+1, 2+2, 4+1, 4+2, 8+2
Number of disks
K+2M, 24, +3 values in between
@@
Table 4: Data load parameters and their value ranges for the storage SSD pool data set; for sequential data loads
Parameter
Value range
Load type
sequential
Block size
128, 256, 512, 1024 KB
Read fraction
0%
Number of jobs
1 - 20
Queue depth
1 - 32
RAID (K+M)
1+1, 2+1, 2+2, 4+1, 4+2, 8+2
Number of disks
K+2M, 24, +3 values in between
@@
Table 5: Data load parameters and their value ranges for the storage HDD pool data set; for sequential data loads
Parameter
Value range
Load type
sequential
Block size
128, 256, 512, 1024 KB
Read fraction
0, 100%
Number of jobs
1 - 20
Queue depth
1 - 32
RAID (K+M)
1+1, 2+1, 2+2, 4+1, 4+2, 8+2
Number of disks
K+2M, 24, +3 values in between
@@
For this study, we collected four data sets for the storage pools and cache under different data loads. The first one is for the cache under the random data loads. To collect it, we generated 510 different data loads using a performance analysis tool Perf [15] in Linux. Table 2 provides the list of the load parameters and the ranges of their values. We quasi-randomly selected these values for each individual data load using the Sobol sequence [16]. This method covers the parameter space more uniformly than random or grid sampling. Each load lasted 60 seconds, and for every second we measured IOPS and average latency for read and write operations separately. In the results, we collected $510\times 60\times 2$ pairs of measurements $(x_{i},y_{i})$.
@@
Similarly, we collected data sets for the SSD pool under random loads. The list of all parameters and their value ranges is in Table 3. In this case, we also used the pool configuration parameters in addition to the data load ones. The first is a RAID scheme, which is defined by the number of data (K) and parity (M) blocks. The second is the total number of disks in the pool. We generated 512 different loads in total for various pool configurations with a duration of 120 seconds each. We measured IOPS and average latency for read and write operations separately every second. In the results, this data set contains $512\times 120\times 2$ pairs of measurements $(x_{i},y_{i})$.
@@
We also collected two data sets for SSD and HDD storage pools under sequential data loads. We used the same procedure described above. The main difference is that only pure loads with read fractions of 0 and 100 % are valuable. One more difference is that larger block sizes are used for the sequential loads. Table 4 and Table 5 show the full list of the data load and configuration parameters and their value ranges. We used the Sobol sequence to generate 512 different loads in total for various pool configurations with a duration of 120 seconds each. We measured IOPS and average latency for read and write operations every second. In the results, each data set contains $512\times 120$ pairs of measurements $(x_{i},y_{i})$.
Models
CatBoost model

### v2

@@ Abstract
High-precision systems modeling is one of the main areas of industrial data analysis. Models of systems, their digital twins, are used to predict their behavior under various conditions. In this study, we developed several models of a storage system using machine learning-based generative models to predict performance metrics such as IOPS and latency. The models achieve prediction errors ranging from 4%–10% for IOPS and 3%–16% for latency and demonstrate high correlation (up to 0.99) with observed data. By leveraging Little’s law for validation, these models provide reliable performance estimates. Our results outperform conventional regression methods, offering a vendor-agnostic approach for simulating data storage system behavior. These findings have significant applications for predictive maintenance, performance optimization, and uncertainty estimation in storage system design.
Index Terms:
Datacenters, digital twin, simulation, storage systems, surrogate modeling.
@@
=-21pt
I Introduction
Data storage systems (DSS) are of vital importance in the modern world. The emergence of big data has necessitated the development of storage systems with a larger capacity and lower costs to efficiently store and process the vast amounts of data generated [2, 3]. Performance modeling is essential in the development of such systems. The modeling allows engineers to analyze the system’s behavior under different conditions and helps them to find the optimal design of the system. It can be used for marketing purposes to estimate the system performance under the customer’s requirements. One more application is diagnostics and predictive maintenance, where model predictions are compared with actual measurements to detect failures and anomalies.
The key components of a typical DSS are controllers, fast cache memory, and storage pools. All data is stored in the pools consisting of several hard disk drives (HDDs) or solid-state drives (SSDs) united with raid schemes. The cache speeds up read and write operations with the most popular data blocks. The controllers are responsible for data management and provide computational resources to process all user input and output requests.
An early attempt to model SSDs was described in [4]. The authors argue that HDD performance models cannot be used for SSD due to certain unique characteristics, such as low latency, slow update, and expensive block-level erasure. The black-box approach to model SSD performance is suggested as it requires minimal a priori information about the device. The researchers found that although black-box simulation does not produce high-quality predictions for HDDs, it can produce accurate ones for SSDs.
@@ I Introduction
Machine learning techniques have proved to be useful in the task of analyzing the disks. As an example, the main approaches discussed in [7] are based on machine learning. The paper uses a methodology that extracts low-cost history-aware features to train SSD performance models that predict response times to requests. The paper also utilizes space-efficient data structures such as exponentially decaying counters to track past activity with $O(1)$ memory and processing cost. Lastly, the paper uses machine learning models such as decision trees, ensemble methods, and feedforward neural networks to predict the storage hardware response time. All of these approaches are designed to make accurate SSD simulators more accessible and easier to use in real-world online scenarios.
@@
Another open-source simulator was developed in [9]. The authors introduce SimpleSSD, a high-fidelity holistic SSD simulator. Aside from modeling a complete storage stack, this simulator decouples SSD parallelism from other flash firmware modules, which helps to achieve a better simulation structure. SimpleSSD processes I/O requests through layers of a flash memory controller. The requests are then serviced by the parallelism allocation layer, which abstracts the physical layout of interconnection buses and flash disks. Although aware of internal parallelism and intrinsic flash latency, the simulator can capture close interactions among the firmware, controller, and architecture.
The idea of SimpleSSD has been further developed in [10]. SimpleSSD 2.0, namely Amber, was introduced to run in a full-system environment. Being an SSD simulation framework, Amber models embedded CPU cores, DRAMs, and various flash technologies by emulating data transfers. Amber applies parallelism-aware readahead and partial data update schemes to mimic the characteristics of real disks.
@@
Although simulating SSD devices might have seemed an exciting research topic, older storage systems have also received the scientific attention they deserve. One of the papers presents a series of HDD simulations to investigate the stiffness of the interface between the head and the disk. In [12], the authors develop a finite-element HDD model using a commercially available three-dimensional modeling package. The results of the numerical analysis provided insight on the characteristics of each type of shock applied to an HDD model.
@@
The evaluation of SSD performance has also been investigated in [15]. SSDs are built using an integrated circuit as memory, so manufacturers are interested in I/O traces of applications running on their disks. This paper proposes a framework for accurately estimating I/O trace execution time on SSDs.
Malladi et al. [16] introduces FlashStorageSim, an SSD architecture performance model validated with an enterprise SSD. It models the SSD controller, flash devices, and host interface (e.g., SATA, PCIe, DDR), allowing for the exploration of various SSD use cases. The model includes mechanisms to improve simulation speed, achieving over 7X reduction in simulation time. Validation against a real SSD shows that FlashStorageSim accurately captures performance trends. The fast-forwarding (FF) and cycle granularity adjustment (CGA) mechanisms further enhance simulation speed. The methodology involves developing FlashStorageSim in C++ to model SSD components and interfaces and incorporating mechanisms to reduce simulation time. The model is validated against a real server SSD using IO benchmarks and trace collection to compare performance metrics like response time and bandwidth. Speed optimization is assessed through fast-forwarding and cycle granularity adjustment, showing simulation speed improvements
The work done by Lebre et al. [17] presents advancements in the simulation of storage resources for distributed computing systems. The main results include the successful extension of SimGrid with comprehensive storage simulation capacities, introducing a versatile API that handles storage components and their contents. The authors characterized the performance of various types of disks, deriving accurate models that account for sharing and contention effects. These models were validated through extensive experiments on the Grid’5000 testbed, demonstrating their practicality and accuracy. The paper also highlighted several potential use cases for the enhanced SimGrid toolkit, such as optimizing data placement and replication policies in MapReduce applications, improving storage management in grids and clouds, and accurately simulating hierarchical storage systems in data centers. The proposed extension thus enables more accurate and comprehensive simulation studies, facilitating better design and management of storage infrastructure in large-scale distributed computing environments.
Louis et al. [18] extends the CloudSim toolkit [19] to include a scalable module for modeling and simulating energy-aware storage in cloud systems. The main results demonstrate that CloudSimDisk effectively simulates the energy consumption of storage systems and validates against analytical models. CloudSimDisk introduces HDD models and power models, supports energy-aware simulations, and includes features for disk array management and transaction processing. Simulations using CloudSimDisk align closely with analytical predictions, confirming its accuracy and efficacy.
The paper [20] by Lee et al. introduces an accurate data-driven model for SSD firmware validation. By directly executing the firmware on both a simulator and the target device, the model collects performance profiles to predict SSD performance. Using a deep neural network (DNN), the model achieves a 3.1% error rate, significantly lower than the 18.9% error from existing simulators. This approach reduces the need for extensive testing on prototype boards, improving efficiency and accuracy in SSD development. This work is the first to apply a data-driven method for SSD performance prediction, effectively bridging the gap between simulation and real-world performance.
[21] Wan et al. develop a simulation model to predict the cumulative distribution function (CDF) of I/O request response time in a RAID 0 system. The methodology involves a bottom-up approach, starting with a single disk drive modeled as an M/G/1 queue. This model is then extended to a RAID 0 system using a split-merge queueing network. The model’s accuracy is validated by comparing simulation results of I/O request response times with measurements from actual devices, showing close agreement. Speed optimizations include fast-forwarding and cycle granularity adjustments, improving simulation efficiency.
This work [22] presents a fork-join queueing simulation model of RAID systems with zoned disk drives, specifically focusing on RAID levels 01 and 5. Starting with an M/G/1 queue-based simulation for a single disk, the model is extended to RAID using a fork-join network. The simulator accepts heterogeneous I/O workloads and produces the I/O request response time distribution, incorporating request-reordering optimizations and validating against device measurements. The simulation model accurately predicts I/O request response times for RAID 01 and 5 systems, showing agreement with actual device measurements. It highlights the impact of request size, workload burstiness, and request-reordering optimizations on performance. The model simulates RAID systems under various conditions, providing insights for improving analytical models and optimizing RAID configurations.
Other work in the context of storage system modeling includes various approaches. Low-level simulators for magnetic tapes [23, 24], hard drives [25], and SSDs [26] use detailed discrete-event simulations. DiskSim [25], which models block-level operations, faces scalability issues. SIMCAN [27] extends this with layered simulations of disks, volumes, and file systems, which can lead to configuration challenges. In HPC simulations, scalability is often prioritized over accuracy, with random I/O times used to improve performance [28].
MapReduce simulators [29, 30] typically use simplistic disk models, focusing on the logic of file systems rather than their performance. This paper extends SimGrid with versatile and accurate storage simulation capabilities, leveraging its expertise in network modeling [31] and introducing an API that addresses sharing and contention effects across different infrastructures.
@@
All of these solutions have a range of limitations. The first one is that they are focused mostly on the modeling of one single device such as HDD or SSD disks. The second is that most of them are based on simulation of physical processes and the firmware stack inside the devices, which require more resources for development. However, one of the main disadvantages is that the models do not take into account vendor specifics of the devices due to the trade secrets. As a result, the errors in performance simulation are quite high for practical use. Finally, HDDs and SSDs are part of DSS with unique architectures and software that influence the device performance and are not taken into account in the solutions described above.
@@
In this work, we present a data-driven approach for DSS modeling that does not have these limitations. All vendor specifics, architecture features, and software impact are learned directly from real system performance measurements. We present the performance modeling and analysis of a data storage cache, SSD, and HDD pools under random and sequential data loads. We provide novel data sets of IOPS and latency measurements for the system’s components for various configurations and load parameters. We consider two different approaches to learn distributions of the performance values, based on parametric and nonparametric generative models. The inclusion of generative models allows for a more comprehensive simulation of system behavior. More precisely, the learned distributions encompass every potential performance value corresponding to the provided data load parameters. By employing these values, we can determine the means and standard deviations of IOPS and latency, and examine the correlations between them. In contrast to conventional regression models, which are limited to predicting only mean values.
We analyze the quality of the model predictions and compare them with a naive baseline algorithm. We also demonstrate a physics-inspired method for checking the reliability of the model predictions. The data sets provided in this work can be used for benchmarking regression algorithms, conditional generative models, and uncertainty estimation methods in machine learning.
Additionally, our approach comprehensively models the entire storage system, encompassing individual disks of various types, the complete array, and the cache layer, with an emphasis on delivering an integrated and holistic solution. This methodology ensures that every component, from the smallest disk to the most complex array configuration, is accounted for in our model. Moreover, our approach is designed to be universally applicable, transcending vendor-specific constraints, thereby ensuring compatibility with a wide array of hardware platforms. This vendor-agnostic nature of our model enhances its versatility, making it suitable for diverse hardware environments and enabling seamless integration across different storage infrastructures.
The article is organized as follows. In Section 2, we define the task of predicting IOPS and latency for storage components under varying workloads and configurations. Section 3 describes the generation of datasets using tools like Perf, covering parameters such as load type, block size, RAID schemes, and job configurations. In Section 4, we introduce CatBoost and Normalizing Flows (NF) for learning performance distributions, alongside a kNN baseline. Section 5 outlines evaluation metrics to assess prediction accuracy and distributional fidelity. Section 6 presents model performance, showing CatBoost’s lower prediction errors (4–10% for IOPS, 3–16% for latency) and NF’s ability to capture complex variability, validated by Little’s law. Section 7 analyzes model behavior, including conservatism in extrapolation regimes, and highlights practical implications for DSS simulation. Finally, Section 8 summarizes findings, advocates for generative models in DSS performance modeling, and proposes future work on uncertainty estimation and explainability.
II Problem statement
TABLE I: Input and output features of the simulation models
Model Inputs Outputs
@@ II Problem statement
Storage pool Load type, IO type, read fraction, block size, number of jobs, queue depth, number of disks, number of data, and parity blocks in RAID IOPS, Latency
@@
Storage cache Load type, IO type, read fraction, block size, number of jobs, queue depth IOPS, Latency
@@
In this study, we consider the simulation of HDD and SSD storage pools and cache, which are the key parts of data storage systems. The goal of this work is to predict the performance of these components for a given configuration and data load parameters. We describe the performance by the number of input and output operations per second (IOPS) and the average of their latencies. The data load is parameterized by: load type, IO type, read fraction, block size, number of jobs, and queue depth for each of these jobs. There are two load types we consider in the work: random and sequential. Each data load consists of a mix of read and write operations. The IO type indicates whether we consider read or write operations in our simulations. The read fraction is the ratio of the number of read operations to the total number of all operations in the data load. Each operation processes one data block of a given size. The data loads are generated by several jobs with a given queue depth.
@@
HDD and SSD storage pools also have configuration parameters. They are the total number of disks in a pool and the software RAID scheme, which is described by the number of data and parity blocks. In this work, the cache has only one configuration and we do not describe it by any parameter. The lists of input and output features in our simulation models are shown in Table I.
@@
Let us denote the vector of input features as $x$, and the vector of outputs as $y$. Also suppose we have $n$ pairs of measurements $\{x_{i},y_{i}\}_{i=1}^{n}$ of output vectors $y_{i}$ for the given inputs $x_{i}$. These pairs are obtained by generating various data loads on a real data storage system and measuring the performance of its pool or cache in terms of IOPS and latency. The goal of this work is to estimate the conditional probability distribution $p(y|x)$ for the storage component using machine learning. Then, we predict the performance $\hat{y}_{j}$ for unknown inputs $x_{j}$ by sampling from the learned distribution:
@@ III Data collection
TABLE II: Data load parameters and their value ranges for the storage cache data set
Parameter Value range
Load type random
Block size 4, 8, 16, 32, 64, 128, 256 KB
Read fraction 0 - 100%
Number of jobs 1 - 64
Queue depth 1 - 16
@@
TABLE III: Data load parameters and their value ranges for the storage SSD pool data set; random data loads
Parameter Value range
Load type random
Block size 4, 8, 16, 32, 64 KB
Read fraction 0 - 100%
Number of jobs 1 - 32
Queue depth 1 - 32
RAID (K+M) 1+1, 2+1, 2+2, 4+1, 4+2, 8+2
Number of disks K+2M, 24, +3 values in between
@@
TABLE IV: Data load parameters and their value ranges for the storage SSD pool data set; for sequential data loads
Parameter Value range
Load type sequential
Block size 128, 256, 512, 1024 KB
Read fraction 0%
Number of jobs 1 - 20
Queue depth 1 - 32
RAID (K+M) 1+1, 2+1, 2+2, 4+1, 4+2, 8+2
Number of disks K+2M, 24, +3 values in between
@@
TABLE V: Data load parameters and their value ranges for the storage HDD pool data set; for sequential data loads
Parameter Value range
Load type sequential
Block size 128, 256, 512, 1024 KB
Read fraction 0, 100%
Number of jobs 1 - 20
Queue depth 1 - 32
RAID (K+M) 1+1, 2+1, 2+2, 4+1, 4+2, 8+2
Number of disks K+2M, 24, +3 values in between
@@
For this study, we collected four data sets for the storage pools and cache under different data loads. The first one is for the cache under the random data loads. To collect it, we generated 510 different data loads using a performance analysis tool Perf [32] in Linux. The tool generated a stream of data blocks from a client to the system for read and write operations and collected statistics about the processing of these blocks. Table II provides the list of the load parameters and the ranges of their values. We quasi-randomly selected these values for each data load using the Sobol sequence [33]. This method covers the parameter space more uniformly than random or grid sampling. Each load lasted 60 seconds, yielding every second IOPS and average latency for read and write operations separately, giving us two separate results with respect to read and write operations. In this case, IOPS means how many data blocks from the client were read or written to the cache per one second. As a result, we collected $510\times 60\times 2=61200$ pairs of measurements $(x_{i},y_{i})$.
@@
Similarly, we collected data sets for the SSD pool under random loads. The list of all parameters and their value ranges is in Table III. In this case, we also used the pool configuration parameters in addition to the data load ones. The first is a software RAID scheme, which is defined by the number of data (K) and parity (M) blocks. The second is the total number of disks in the pool. The minimum amount of disk in the pool can be determined by applying the K+2M formula listed in the table. E.g. for RAID schema 4+2, the minimum number of disks required is 8, and the maximum is 24 for any configuration. We include minimum, maximum, and 3 values in between, like 12, 16, and 19 disks in that example. We generated 512 different loads in total for various pool configurations with a duration of 120 seconds each. We measured IOPS and average latency for read and write operations separately every second. In the results, this data set contains $512\times 120\times 2=122820$ pairs of measurements $(x_{i},y_{i})$.
@@
We also collected two data sets for SSD and HDD storage pools under sequential data loads. We used the same procedure described above. The main difference is that only pure loads with read fractions of 0 and 100 % are valuable. One more difference is that larger block sizes are used for sequential loads. Table IV and Table V show the full list of the data load and configuration parameters and their value ranges. We used the Sobol sequence to generate 512 different loads in total for various pool configurations with a duration of 120 seconds each. We measured IOPS and average latency for read and write operations every second. In the results, each data set contains $512\times 120=61440$ pairs of measurements $(x_{i},y_{i})$. As a result, we collected 4 distinct data sets, with different amounts of rows. Additionally, we engendered one feature by multiplying the depth and number of jobs and splitting the raid schema variable into two variables. Instead of string feature K+M therefore obtaining K and M. Thus, for SSD and HDD pools modeling we use 10 features, and for cache - only 7.
The value ranges were derived from the system specification. Due to the sheer size of the parameter ranges, and the amount of all possible combinations, it is more sensible to use only a subsample of the vast parameter space. The Sobol sequence was chosen because it ensures more uniform coverage of the parameter space compared to random selection. It is important because it ensures to sample from rare or critical regions, and that samples are equally distanced to avoid clustering in high-dimensional space. Thus, leading towards better generalization ability of the model.
IV Models
IV-A CatBoost model